3. The probability that a bulb produced by a factory will fuse after certain period of time is 0.05. Find the probability that out of 5 such bulbs
a. None fuses
b. Not more than 2 fuse
c. More than 2 fuse

Answer by stanbon(75887)

The probability that a bulb produced by a factory will fuse after certain period of time is 0.05. Find the probability that out of 5 such bulbs
a. P(None fuses) = 0.95^5 = 0.7738
----------------------------------------------
b. P(Not more than 2 fuse) = binomcdf(5,0.05,2) = 0.9988
----------------------------------------------
c. P(More than 2 fuse) = 1-0.9988 = 0.0012
==============================================
Cheers,
Stan H.
================